What are the slopes of the following binomials
1. 4a+4b-8=4c-2a
2. 3x+4y+-12
3. -9=3y-x
4.24=-2x-6y
5.48=12y-24x

Answer by Alan3354(69443)

What are the slopes of the following binomials
1. 4a+4b-8=4c-2a
2. 3x+4y+-12
3. -9=3y-x
4.24=-2x-6y
5.48=12y-24x
----------------
Equations on an x-y plane have a slope.
1 has a, b & c as variables.
2 has no equal sign --> not an equation.
3. -9=3y-x
Solve for y:
y = (1/3)x - 3
The coefficient of x is the slope, m.
m = 1/3
============
24=-2x-6y
Solve for y:
The coefficient of x is the slope, m.
---
48=12y-24x
Same as #4
==============
PS None of these are binomials.